### v3

# RGB-Th-Bench: A Dense benchmark for Visual-Thermal Understanding of Vision Language Models

Abstract
We introduce RGB-Th-Bench, the first benchmark designed to evaluate the ability of Vision-Language Models (VLMs) to comprehend RGB-Thermal image pairs. While VLMs have demonstrated remarkable progress in visual reasoning and multimodal understanding, their evaluation has been predominantly limited to RGB-based benchmarks, leaving a critical gap in assessing their capabilities in infrared vision tasks. Existing visible-infrared datasets are either task-specific or lack high-quality annotations necessary for rigorous model evaluation. To address these limitations, RGB-Th-Bench provides a comprehensive evaluation framework covering 14 distinct skill dimensions, with a total of 1,600+ expert-annotated Yes/No questions. The benchmark employs two accuracy metrics: a standard question-level accuracy and a stricter skill-level accuracy, which evaluates model robustness across multiple questions within each skill dimension. This design ensures a thorough assessment of model performance, including resilience to adversarial and hallucinated responses. We conduct extensive evaluations on 19 state-of-the-art VLMs, revealing significant performance gaps in RGB-Thermal understanding. Our results show that even the strongest models struggle with thermal image comprehension, with performance heavily constrained by their RGB-based capabilities. Additionally, the lack of large-scale application-specific and expert-annotated thermal-caption-pair datasets in pre-training is an important reason of the observed performance gap. RGB-Th-Bench highlights the urgent need for further advancements in multimodal learning to bridge the gap between visible and thermal image understanding. The dataset is available through this link, and the evaluation code will also be made publicly available.
1 Introduction
Figure 1: Two data samples from RGB-Th-Bench. Due to space limit, here we show only a subset of 16 Q/As across 4 skill dimensions for each sample, and the response from 3 selected VLMs.
Vision-Language Models (VLMs) [8, 52, 53, 12, 57, 39, 38, 32, 6, 47, 3, 44] have made remarkable progress in recent years, leveraging large-scale image-text datasets to align visual representations with the powerful reasoning capabilities of Large Language Models (LLMs) [43, 49, 17, 7, 11]. To systematically evaluate their multimodal understanding, numerous benchmarks have been developed, such as MME [18], MMBench [40], SEED-Bench [27], and MM-Vet [54]. These benchmarks have become widely recognized for assessing various vision-language skill dimensions, including core perception tasks, domain knowledge, reasoning abilities, and application-specific tasks.
Despite the success of existing VLM benchmarks, they primarily focus on RGB images and lack comprehensive evaluation for multimodal tasks involving infrared (thermal) imagery. Visible-infrared paired datasets are rare and often designed for specific applications, such as pedestrian detection, autonomous driving, military applications, or object tracking [15, 10, 20, 9, 5, 23, 22, 24, 4]. Moreover, a critical gap is the absence of high-quality, expert-annotated datasets with aligned visible-thermal captions, which are essential for training and evaluating VLMs on thermal image understanding. As a result, the evaluation of existing VLMs on RGB-Thermal paired data is an unexplored area in vision-language research.
To this end, we introduce RGB-Th-Bench, the first benchmark designed to rigorously evaluate the ability of VLMs to comprehend RGB-Thermal image pairs.
RGB-Th-Bench currently comprises over 1,600 Yes/No questions across 14 skill dimensions. Each skill dimension includes four questions per image, resulting in 116 questions per skill dimension and 56 questions per RGB image in total. This design ensures that RGB-Th-Bench is both dense and comprehensive.
The benchmark employs two accuracy metrics: a question-level accuracy metric, where random performance is 50%, and a stricter skill-level accuracy metric, where all four questions within a skill dimension must be answered correctly for a ”pass”. This dual-metric approach allows for a nuanced evaluation of model’s performance including its robustness against hallucinations and adversarial questions.
To evaluate RGB-Th-Bench, we conduct extensive experiments on a diverse set of 19 VLMs, including both open-source and closed-source models, reporting their performance across various skill dimensions and metrics. Our results reveal significant performance gaps in RGB-Thermal understanding. Notably, we find that most models struggle with RGB-Thermal tasks, and their performance is ultimately upper-bounded by their core RGB-based capabilities.
Our benchmark highlights the need for substantial improvements in VLMs before they can be effectively deployed in real-world RGB-Thermal applications. We hope RGB-Th-Bench serves as a foundational resource for advancing multimodal research and improving vision-language models in the challenging domain of thermal image understanding.
Inspired by ”Machine Learning that Matters” [50], we believe that designing impactful benchmarks and metrics can drive upstream research efforts, guiding the selection of training datasets, structuring of experiments, and definition of objective functions. To this end, this work demonstrates how improvements in VLM accuracy can translate into meaningful impact for real-world problem domains.
2 Related Work
Dataset #Thermal RGB-Th Pairs Text Ann. Task Obj. Ctg. Features
[…]
CVC-13 [10] 23 Yes No Ped Det Peds Small size, urban focus
CVC-14 [20] N/A Yes No Ped Det Peds Misaligned pairs, no fusion
CVC-15 [9] 100 Yes No Ped Det Peds BB ann. provided
KAIST [22] 95k+ Yes No Self-Driving, Ped Det Ped, Veh Well-aligned, day/night
LLVIP [24] 15,488 Yes No LL Ped Det Peds High-quality alignment
INO [23] N/A (video) Yes No Obj Det, Ped Track Peds Weather-varied, video
FLIR [4] 26,442 No No Self-Driving, Obj Det 15 Ctg. BB ann. provided
TNO [5] 63 Yes No Img Fus Mil objs Multispectral (Vis, NIR, LWIR)
27UAV [19] 27 Yes No UAV Surv N/A Small scale, UAV focus
Table 1: Comparison of thermal image datasets. Abbreviations: #Thermal = Number of thermal images; RGB-T Pairs = Paired RGB-thermal images; Text Ann. = Text or caption annotations; Task = Task focus (Det: Detection, Ped: Pedestrian, Wpn: Weapon, $L$: Low-light, Obj: Object, Track: Tracking, Img Fus: Image Fusion, UAV Surv: UAV-based Surveillance); Obj. Ctg. = Object categories (Veh: Vehicles, Mil objs: Military objects, 15 Ctg.: 15 annotated categories); Features: Unique dataset features ($B$ ann.: Bounding box annotations, Vis, NIR, LWIR: Visible, near-infrared, longwave-infrared data).
Vision-Language Models. Building on the remarkable success of LLMs [43, 49, 17, 7, 11], recent research has focused on VLMs [8, 52, 53, 12, 57, 39, 38, 32, 6, 47] to enhance multimodal understanding. These models achieve this by aligning the visual features extracted from pre-trained image encoders with LLMs using large-scale image-text datasets.
Recent Benchmarks for Vision-Language Models. To systematically assess the capabilities of VLMs, several benchmarks have been introduced, such as MME [18], MMBench [40], SEED-Bench [27], MMMU [56], MM-Vet [54], and NaturalBench [29]. These benchmarks have become widely recognized as standard evaluation tools, as reflected in their prominence on performance leaderboards [14].
Despite their broad adoption, these benchmarks primarily assess single-image text comprehension and lack support for tasks involving multiple images or interleaved image-text sequences. While this was a reasonable limitation in earlier VLMs designed for single-image inputs, recent advancements [57, 52, 53, 43, 44] have enabled models to process multiple images, sequential video frames, and arbitrarily interleaved multimodal inputs. In response to this shift, new benchmarks such as MM-Vet2 [55], SeedBench2 [26], and Seed-Bench2-Plus [28] have been introduced as extensions to their predecessors.
Most VLM benchmarks evaluate model performance across multiple skill dimensions, which can be broadly categorized into four overlapping domains:
  • Core perception skills, including scene understanding, object existence, object attributes, and spatial reasoning [18, 54, 27, 40, 34].
  • Domain knowledge, encompassing both general knowledge and scientific understanding [41, 56, 18, 26].
  • Reasoning abilities, including mathematical reasoning, logical reasoning, commonsense reasoning, and attribute-based reasoning [18, 56, 25, 51].
  • Application-specific skills, such as celebrity recognition [18, 40], meme comprehension [26], landmark recognition [26], chart interpretation [28], map reading [28], OCR [54], GUI navigation [35], and screenshot-to-code translation [46].
Visible-Infrared Image Pair Datasets. Infrared image datasets are considerably rarer than their visible-light counterparts, with most existing datasets designed for specific tasks such as pedestrian detection, object tracking, and image fusion. Table 1 presents the comparison among these datasets. Notably, datasets like OTCBVS [15], and and FLIR [4] only provide Thermal images, which limits their use for multimodal learning. CVC-13/14/15 [10, 20, 9], 27UAV [19], and TNO [5] are small in scale, containing less than 100 RGB-Thermal pairs of highly similar scenes. Although KAIST [22], and LLVIP [24] provide larger collections, they remain task-specific with limited object diversity, limiting their broader applicability in vision-language research.
Despite these efforts, no existing dataset provides captioned infrared images or visual instruction annotations, which are essential for advancing multimodal understanding of thermal imagery. Additionally, key applications such as heat and energy loss detection remain underrepresented due to annotation challenges, as infrared images often lack well-defined object boundaries. The absence of such datasets highlights the need for a more comprehensive benchmark that integrates textual descriptions with aligned visible-infrared image pairs, addressing a critical gap in vision-language model evaluation.
3 RGB-Thermal Image Benchmark
In this work, we developed a high-quality dense multimodal benchmark for visual-thermal understanding, called RGB-Th-Bench, that requires comprehensive skills or capabilities to stimulate the research and provide a foundation for future advancements in infrared image analysis and visual-thermal multimodal research.
3.1 Data Collection
The benchmark includes a total of 1624 questions and 58 images (29 RGB-Thermal sample pairs). Similar to MME [18], POPE [34], and NatrualBench [29], these questions entail Yes/No responses. We do not use any image or question from any other dataset or benchmark. The images are either manually collected from internal documents or directly taken using ”FLIR ONE Edge Pro” camera [1]. All the questions and their ground truths answers are manually annotated by us, under the supervision of domain experts. As shown in Figure 2 and Figure 3, most of the scenes are of importance for risk and anomaly detection in industrial or residential settings for example for detection of poor insulation, gaps and cracks, leakage, overheating, electrical issue, and fire. Intentionally, this benchmark is not designed for military or surveillance applications.
Figure 2: Some RGB-thermal pair samples in residential settings
Figure 3: Some RGB-thermal pair samples in industrial settings
3.2 Evaluation Dimensions
For each sample pair, there are a total of 56 questions across 14 core evaluation or skill dimensions, resulting in four independent Yes/No questions per skill dimension for each sample. This way, we make sure that every sample pair is thoroughly tested from all possible skill dimensions to make our benchmark a dense evaluation setting, which is unlike the benchmarks that ask only one or two questions from each sample image, like MME [18], MM-Vet [54] and Seed-Bench2 [26]. These skills are grouped into two types of prompt:
  • Single RGB & Text (RGB-Txt): This prompt-group is necessary to evaluate model performance without thermal input, so that we can control the effect of addition of the thermal image to the same input. This prompt-group consists of seven diverse evaluation dimensions including Scene Understanding (Scene), Detailed Object Presence (ObjPr), Instance Attributes (Attr), Instance Location w.r.t. Image (LocIm), Instance Spatial Relation (SpRel), Instance Counting (Count), and Instance Interaction (Interact).
  • RGB-Thermal pair & Text (RGB-Th-Txt): This prompt-group is to evaluate model performance with thermal input, and similarly includes seven diverse evaluation dimensions, including Temperature to Color Mapping Understanding (Temp2Col), RGB-Thermal-Heatmap Alignment (HeatAlign), Instance Thermal Attribute (ThermAttr), Relative Thermal Attribute (RelTherm), Warmest Areas Detection (Warmest), Coldest Areas Detection (Coldest), and Risk and Anomaly Detection (Anomaly).
The detailed guideline for annotations of all evaluation dimensions is provided in the supplementary materials.
Benchmarks #Skills Visual Modality Visual Source #Q/A Annotations #Q/A Source Answer Type Question Density
MMBench [40] 20 RGB(s) mixed 3217 human free-form 1
HallusionBench [21] 9 RGB web 1129 human Y/N 3.26
MMVet [54] 6 RGB mixed 218 human free-form 1.09
POPE [34] 1 RGB MSCOCO [36] 3000 mixed Y/N 6
SEEDBench2 [26] 27 RGB(s) & video other datasets 24371 human&GPT A/B/C/D 1
LLaVA-Bench [39] 3 RGB web 64 human free-form 2.5
MME [18] 14 RGB mixed 2374 human Y/N 2
RGB-Th-Bench(Ours) 14 RGB-Thermal self-sourced 1624 human Y/N 56
Table 2: Comparison of RGB-Th-Bench with some other established benchmarks. ”#Skills” and ”Question Density” mean the number of skill dimensions and the number of questions per each visual sample, respectively.
3.3 Instruction Design
To enable the collection of quantitative performance statistics, the instruction design is structured to elicit ”Yes” or ”No” responses from the model. Consequently, the instruction comprises two components: a concise question and a fixed context. The context for all questions in the ”RGB-Txt” prompt-group is Based on this image, answer the following question with strictly either ”Yes” or ”No”, without any extra explanation, and the one for questions in the ”RGB-Th-Txt” prompt-group is Based on these two images, and the fact that the second image is the thermal image taken from the same scene as the first image, answer the following question with strictly either ”Yes” or ”No”, without any extra explanation.
If the model’s response does not include Yes or No (case insensitive), it is marked as a disregard to the simple Yes/No instruction and considered as a “Fail”. Note that we do not require strict exact matching, for example the model can respond like ”The answer is yes” and we qualify it as a ”Yes”, since it includes either case-insensitive ”Yes” or ”No”. Similar to MME [18], we argue that a good VLM (especially after instruction tuning) should be able to follow such a simple instruction, which is also very common in everyday life.
3.4 Evaluation Strategy
Since the output of the model is limited to two types (“Yes” or “No”), it is convenient to measure the accuracy metric on both question-level and skill-level for each skill dimension:
  • Question-level Accuracy (QAcc): The percentage of the questions answered correctly, with random baseline score of 50%.
  • Skill-level Accuracy (SAcc): if the model correctly answers all the four questions of a skill dimension of a sample pair, then we consider that as a “Pass” for that skill, otherwise a “Fail”. The random baseline score is 6.25% in this case.
It can be seen that Skill-level Accuracy is a stricter measurement but also better reflects the comprehensive understanding degree of the model in its respective skill. In addition, we calculate the overall score of each prompt-group based on the average of accuracy score in their skill dimension.
3.5 Advantages of RGB-Th-Bench
Table 2 provides a comparison between our work and some other established VLM benchmarks.
RGB-Th-Bench is the first and the only benchmark that focuses on both RGB and thermal understanding, and also introduces several key advantages that make it a robust tool for advancing research in visual-thermal understanding and infrared image analysis. Below, we outline the primary benefits:
  • Flexibility and Robust Metrics: RGB-Th-Bench provides flexibility in evaluation by allowing for the inclusion of multiple questions per evaluation dimension. If the scene lacks sufficient number of objects to generate diverse questions, multiple questions can focus on specific objects, often incorporating challenging or adversarial cases, such as negative questions. This ensures a more robust assessment of a model’s resilience to adversarial inputs and susceptibility to hallucination.
  • Fairness in Evaluation: Unlike multiple-choice question-answering benchmarks, where the model benefits from having all answer options within its context and knowing implicitly that only one option is correct, RGB-Th-Bench uses independent questions. Each query stands alone, requiring the model to analyze and respond without reliance on contextual cues from a predefined set of options. This independence ensures a more equitable and unbiased evaluation of the model’s capabilities.
  • Dense and Comprehensive Testing: RGB-Th-Bench rigorously evaluates models across 14 core aspects of visual-thermal understanding by carefully crafting 56 questions for each RGB-Thermal pair sample. Instead of posing a single or limited set of questions per image, our benchmark tests the model comprehensively, examining its robustness and generalization across multiple dimensions. This dense approach ensures a thorough assessment of the model’s strengths and weaknesses, promoting a deeper understanding of its performance.
  • High-Quality Annotations: All questions and answers in RGB-Th-Bench are manually annotated by us with the supervision of domain experts, ensuring their relevance and accuracy. Additionally, the images used in the benchmark are sourced exclusively from internal documents or directly taken by us, and not from other datasets or from the web, guaranteeing that they are not included in pre-training or fine-tuning datasets. This ensures a true test of the model’s generalization capability and eliminates potential biases arising from prior exposure to the data.
4 Experiments
4.1 Target Models
We evaluate a total of 19 recent VLMs that support multi-image inputs, including Chameleon-7B [47], InternVL2.5 series [13], SmolVLM [42], Llava-Next series [37, 30, 33], Pixtral-12B [6], Llava-Onevision-ov series [31], MiniCPM-V-2.6 [53], MiniCPM-o-2.6 [2], Qwen2-VL [52] and Qwen2.5-VL [48] series, and GPT4o [43]. To facilitate the evaluation process, we incorporate VLMEvalKit [16] as our evaluation toolkit.
Models Language Model Vision Model RGB-Txt RGB-Th-Txt Overall
QAcc SAcc QAcc SAcc QAcc SAcc
Random Baseline - - 50 6.25 50 6.25 50 6.25
Statistically Significant - - >>>53 >>>8 >>>53 >>>8 >>>52 >>>7.5
Chameleon-7B [47] Unified-Mixed-Modal Unified-Mixed-Modal 26.48 0.49 31.54 0.49 29.01 0.49
InternVL2.5-1B [13] Qwen2.5-0.5B InternVit-300M-v2.5 65.89 23.15 56.68 4.93 61.28 14.04
InternVL2.5-2B [13] InternLM2.5-1.8B InternVit-300M-v2.5 67.86 25.62 61.60 5.42 64.73 15.52
Llava-Onevision-1B-ov [31] Qwen2-0.5B SigLip-400M 69.58 30.05 47.66 4.93 58.62 17.49
SmolVLM-2.3B [42] SmolLM2-1.7B SigLip-400M 69.09 28.57 53.45 9.85 61.27 19.21
Qwen2-VL-2B-Instruct [52] Qwen2-1.5B QwenViT 75.25 35.96 58.64 4.93 66.95 20.44
Llava-Next-Mistral-7.6B [37] Mistral-7B CLIP-ViT-L/14 77.83 39.41 52.82 3.45 65.33 21.43
Llava-Next-Interleave-8B [33] Qwen1.5-7B SigLip-400M 76.72 40.39 47.30 4.93 62.01 22.66
Llava-Next-Llama3-8B [30] Llama3-8B-Instruct CLIP-ViT-L/14 77.22 38.42 58.63 10.34 67.92 24.38
Pixtral-12.4B-2409 [6] Mistral-Nemo-12B Pixtral-ViT-400M 76.85 39.41 59.13 9.36 67.99 24.38
Qwen2.5-VL-3B-Instruct [48] Qwen2.5-7B QwenViT 80.17 46.80 53.68 5.42 66.93 26.11
Llava-Onevision-8B-ov [31] Qwen2-7B SigLip-400M 79.19 42.86 60.73 10.84 69.96 26.85
InternVL2.5-4B [13] Qwen2.5-3B InternViT-300M-v2.5 79.31 43.84 59.13 9.85 69.22 26.85
InternVL2.5-8B [13] InternLM2.5-7B InternViT-300M-v2.5 79.68 46.31 63.08 8.37 71.38 27.34
Qwen2.5-VL-7B-Instruct [48] Qwen2.5-7B QwenViT 75.62 37.44 67.26 17.24 71.44 27.34
MiniCPM-V-2.6-8B [53] Qwen2-7B SigLip-400M 79.80 45.32 63.31 9.85 71.56 27.59
MiniCPM-o-2.6-8B [2] Qwen2.5-7B SigLip-400M 81.16 46.31 62.20 9.85 71.68 28.08
Qwen2-VL-7B-Instruct [52] Qwen2-7B QwenViT 81.16 46.80 60.61 9.85 70.88 28.33
GPT4o-20241120 [43] closed-source closed-source 84.11 56.16 60.85 13.79 72.48 34.98
Table 3: $M$-Vet v2 evaluation results on various VLMs regarding each prompt-group and the overall average score. For each column, the highest figure is shown Bold. Among the open-weight models, the first, the second, and the third highest figures are highlighted by green, yellow and blue backgrounds, respectively. All the numbers are presented in % and the full score is 100%. The conditions in the ”Statistically Significant” row are thresholds for an accuracy score to be statistically significant with p_value<<<0.05 in its respective column.
4.2 Main Results
Table 3 presents the evaluation results of various VLMs on both the RGB-Txt and RGB-Th-Txt prompt groups, along with their overall performance.
In the RGB-Th-Txt prompt group, Qwen2.5-VL-7B-Instruct significantly outperforms other VLMs, including the closed-source GPT4o, achieving a QAcc score of 67.26% and an SAcc score of 17.24%. Additionally, Mini-CPM-V-2.6, InterVL2.5-8B, and Llava-OneVision-8B-ov emerge as top-performing open-weight models.
Conversely, in the RGB-Txt prompt group, GPT4o leads the performance rankings. However, Qwen2-VL-7B-Instruct, MiniCPM-o-2.6, and Qwen2.5-VL-3B-Instruct trail behind by 3% in QAcc and 10% in SAcc. This suggests GPT4o exhibits stronger robustness in this prompt setting. Notably, Qwen2.5-VL-3B-Instruct surpasses many larger models, including Qwen2.5-VL-7B-Instruct, which ranks in the lower half of this prompt group, in contrast to its superior performance in the RGB-Th-Txt setting.
Overall, GPT4o ranks first, with MiniCPM-o-2.6 and Qwen2-VL-7B-Instruct emerging as strong contenders in QAcc and SAcc, respectively. Meanwhile, Chameleon-7B demonstrates the lowest performance across all prompt groups and accuracy metrics, primarily due to its poor instruction-following capability.
To further illustrate model performance across different evaluation dimensions, Figures 4 and 5 provide visual representations of QAcc and SAcc scores, where higher accuracy is indicated in green.
Figure 4: Illustration of models’ QAcc scores across all skill dimensions, with 50% being the random baseline performance, and where green represent higher values.
Figure 5: Illustration of models’ SAcc scores across all skill dimensions, with 6.25% being the random baseline performance, and where green represent higher values.
4.3 Observations
Through an extensive evaluation of VLMs on RGB-Th-Bench, we derive key insights that can inform future research.
The disparity in performance between RGB-Txt and RGB-Th-Txt underscores the increased complexity of comprehending RGB-Thermal data. All VLMs perform worse on the latter prompt, which we attribute to two primary factors:
  • First, success in RGB-Thermal-specific skills depends on strong performance in core RGB-based skills. For instance, as illustrated in Figure 1, answering ”Instance Thermal Attribute” questions requires proficiency in ”Detailed Object Presence” and ”Instance Attributes” (e.g. color). Similarly, ”RGB-Thermal-Heatmap Alignment” questions necessitate strong performance in ”Detailed Object Presence,” ”Instance Attributes” (e.g. color), and ”Instance Location w.r.t. Image” (e.g. shift or zoom). This suggests that RGB-Thermal performance is upper-bounded by RGB-only comprehension skills.
  • Second, as discussed in Section 2, there is currently no large-scale, expert-annotated thermal-caption dataset available for pre-training or fine-tuning. While some thermal-caption samples may exist in general image-caption datasets, their scarcity likely limits model capabilities in thermal understanding.
Performance varies significantly across different evaluation dimensions. This can be attributed to two key reasons:
  • First, the complexity of evaluation dimensions varies substantially. Certain tasks demand more sophisticated reasoning than others. For example, as shown in Figure 1, all three VLMs ”Pass” the SAcc criteria for ”Scene Understanding” and ”Detailed Object Presence,” yet they ”Fail” for ”Instance Spatial Relation” and ”Instance Interaction.” A qualitative comparison reveals that the latter two require deeper contextual comprehension and reasoning.
  • Second, the distribution of training data across different skills plays a crucial role. Large-scale pre-training datasets such as LAION 400M [45] primarily consist of image-caption pairs, which align well with skills like ”Scene Understanding,” ”Detailed Object Presence,” and ”Instance Attributes.” Consequently, models demonstrate stronger performance in these dimensions across both QAcc and SAcc metrics.
SAcc serves as a stricter and more informative metric. For instance, as illustrated in Figure 1, three of the four questions from ”Temperature to Color Mapping Understanding”, which are about a color sequence, are negative questions to each other—meaning a correct ”Yes” response to one question necessitates a ”No” response to the others. However, both GPT4o and MiniCPM-o-2.6 failed in this category, highlighting the value of SAcc in capturing comprehensive understanding.
Most VLMs adhere to the simple ”Yes/No” response format, with a few exceptions. Notably, Chameleon [47], Llava-OneVision-1B-ov [32], and SmolVLM [42] failed to comply with this instruction 712, 6, and 2 times, respectively. Additionally, GPT4o applied content moderation by refusing to answer 28 questions. In all such cases, we classified them as ”Fail.” A comparison with prior benchmarks, such as MME [18] and MMBench [40], which reported significant instruction-following issues, suggests that recent VLMs exhibit improved alignment in this regard.
5 Conclusion
In this work, we introduced RGB-Th-Bench, the first comprehensive visual-thermal evaluation benchmark for Vision-Language Models (VLMs). RGB-Th-Bench is distinguished by its rigorous design, including its distinct purpose, high-quality data source, dense annotation coverage, and strict evaluation metrics. Our benchmark provides a robust and strict metric to assess the capabilities of VLMs in handling RGB-thermal paired data, addressing a critical gap in multimodal research.
We conducted extensive evaluations of various VLMs on RGB-Th-Bench, analyzing their performance to offer insights into different model selections. Our findings reveal that even the best-performing models, Qwen2.5-VL [48] and GPT4o [43], achieve less than 18% on our stringent SAcc metric (random baseline: 6.25%) for RGB-Thermal-Text prompt group. Given the potential of VLMs as multimodal agents, a fundamental requirement is the ability to process diverse internal documents and generate accurate responses. The suboptimal performance observed highlights the significant progress still needed before VLMs can be effectively deployed for RGB-thermal applications.
Despite its contributions, this work has the following limitations:
  • Limited Dataset Size: Acquiring and annotating high-quality thermal data is costly, particularly with the detailed annotations and comprehensive skill sets required in RGB-Th-Bench.
  • Restricted Response Format: The benchmark employs Yes/No responses to simplify accuracy measurement, eliminating the need for prompt engineering or human/LLM-based evaluations. However, this constraint limits expressive open-ended evaluations.
We hope this work provides valuable insights and serves as a foundation to guide future advancements in vision-language multimodal research, particularly in the challenging domain of RGB-thermal understanding.
Supplementary Material
6 Guidelines for Evaluation Dimensions
6.1 Single RGB & Text
  • Scene Understanding (Scene): Each question is about the major content of the image, in a way that one should be able to answer the questions by having a glimpse over the whole image, and not having to directly look at individual objects or people in detail. The questions are not related to individual objects or text in the image, but to the overall theme of the picture.
  • Detailed Object Presence (ObjPr): Each question is about the presence of an object in the scene. The object can be anything with contained shape (chair, car, text) or unlimited (sky, sea, wall, floor), except text or numbers. One should be able to answer the questions by mainly focusing on the presence of the objects, rather than other aspects, such as color, shape, spatial location or its identity.
  • Instance Attributes (Attr): Each question is about some visual attributes of a specific present object such as color, material, shape or fine-grained type. One should be able to answer the questions by looking at the visual appearance of the mentioned object, without considering other aspects, such as spatial location or its identity.
  • Instance Location w.r.t. Image (LocIm): Each question is about location of a specific present object. The location is with respect to the image (top, bottom, left, right, center) and not relative to other objects.
  • Instance Spatial Relation (SpRel): Each question is about the spatial relation between two present objects in the image. The location is relative to each other and in 3D space (up/down , left/right, back/front).
  • Instance Counting (Count): Each question is about the number of appearance of a present object type. The minimum count is one.
  • Instance Interaction (Interact): Each question is about the non-spatial relations and connections between two present objects. One should be able to answer the questions by finding the two mentioned objects, and slightly reason over the image to understand their interaction.
6.2 RGB-Thermal pair & Text
  • Temperature to Color Mapping Understanding (Temp2Col): One or two question is dedicated to the presence of any color to temperature scale bar, and its information like the temperature unit (F or C), temperature range, etc. Other questions are about the mapping of color to temperature. One example of a question form could be “Is the the following sequence of colors correctly represent temperature distribution from low to high: [sequence of colors (at least 3 colors)] ?”
  • RGB-Thermal-Heatmap Alignment (HeatAlign): Questions are either about the difference between the thermal image viewpoint and the RGB image viewpoint, or about the alignment of the heatmap on the black edges of the objects (if the thermal image has an overlay edge filter). Some example forms are:
    • ”Is the thermal image viewpoint exactly matching the RGB image viewpoint without any shift or zoom?”,
    • ”Is the thermal image viewpoint exactly matching the RGB image viewpoint without any shift or zoom?”,
    • ”Is the thermal image slightly shifted to the right/left compared to the RGB image viewpoint?”
    • ”Is the thermal image a zoomed-in crop from the center/top-left/bottom/… viewpoint of the RGB image?”
    • ”Should the heatmap be shifted down/right/left/up or zoom out/in to correctly match the black edges of the objects?”
  • Instance Thermal Attribute (ThermAttr): Each question is about some thermal attributes such as temperature intensity or density of a specific present object. One should be able to answer the questions by finding the object in both RGB and thermal images and carefully look at the visual appearance of the object in the thermal image, but does not have to consider information of other aspects, such as spatial location or its identity.
  • Relative Thermal Attribute (RelTherm): Each questions is about the relative difference between temperature intensity or density of two present objects. One should be able to answer the questions by finding the two objects in both RGB and thermal images and carefully look at the visual appearance of the object in the thermal image, but does not have to consider information of other aspects, such as spatial location or its identity.
  • Warmest Areas Detection (Warmest): Each question is about if a specific present area is the warmest area.
  • Coldest Areas Detection (Coldest): Each question is about if a specific present area is the coldest area.
  • Risk and Anomaly Detection (Anomaly): Each question is about if the thermal state of a specific present object may be considered as a risk or anomaly based on the context of the RGB and thermal images and the nature and purpose of the object. Risks and anomalies are reasonable and may be related to any poor insulation, gaps and cracks, leakage, overheating, electrical issue, fire issue, or any other issue based on the context and the function of the objects.